A man is standing at a basket ball court . The total area of that court is 10sq ft . so what is the perimeter of that ground

Grade:8

3 Answers

Arun
4 years ago
Dear Shivani
 
Please tell is basketball court square shape or any other shape. For any other shpae, you need to given any one information extra
Vikas TU
4 years ago
Dear student 
Basket ball court is square shape so length = sqrt(10) 
Perimeter = 4*l 
Perimeter = 4* sqrt(10) 
Good Luck 
Yashraj
3 years ago
area of basket ball court =l2
10 = l2
l=root 10
perimeter of basket ball court =4*l
perimeter =4*101/2